Sol Summary Report Dec 23rd

Crew 236 Sol Summary Report 23DEC2021

Sol: 4

Summary Title: Groundhog Day

Author’s name: Kasey Hilton

Mission Status: Nominal

Sol Activity Summary: It’s not Groundhog Day but it feels like we lived yesterday all over again. After having to cut yesterday’s EVA short due to radio issues, we retried the EVA today with the same wake up time, same breakfast, same EVA crew, same destination, but without the radio failure. Meanwhile, the remaining crew at the Hab spent time catching up on research and a few good books. After reuniting over bowls of chicken noodle soup, we continued with research and individual free time. The plan is to watch Airplane on someone’s small laptop screen to top the night off. Too bad the Hab doesn’t have any Bill Murray movies.

Look Ahead Plan: The forecast for tomorrow includes some rain, so we will request an EVA and see if the weather allows for EVA tomorrow. If no EVA is possible, the crew will spend a day inside working on research, outreach and free time.

Anomalies in work: None

Weather: Partly overcast, temperature in the 40s

Crew Physical Status: Nominal

EVA: Today’s EVA was successful with both research objectives achieved. A crew member’s spacesuit fan died towards the end, but the crew quickly returned to the Hab after.

Reports to be filed: Operations Report, GreenHab Report, EVA Report, Health and Safety Report, Sol Summary, and Journalist Report (with photos).

Support Requested: It is very windy tonight so is there anything special we need to do to make sure doors don’t blow open?

Operations Report Dec 23rd

Crew 236 Operations Report 23DEC2021

SOL: 4

Name of person filing report: Pavithra Ravi

Non-nominal systems: None

Notes on non-nominal systems: N/A

Generator:

Hours run: 16 hours 2 minutes

From what time last night: 16:58

To what time this morning: 09:00

List any additional daytime hours when the generator was run: None

Solar— SOC % (Before generator is run at night): Information provided by Outpost

Station Propane Reading – % Information provided by Outpost

Ethanol Free Gasoline: Not in use

Water (loft tank): 32.5 gallons

Water Meter: 1539893 units

Water (static tank): approx. 370 gallons

Static to Loft Pump used: No

Water in GreenHab: approx. 213 gallons

Water in ScienceDome: N/A

Toilet tank emptied: No

Deimos rover used: No

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Spirit rover used: No

Hours (before EVA 1): N/A

Beginning charge (Before EVA 1): N/A

Ending charge (On return from EVA 1, before recharging): N/A

Hours (before EVA 2): N/A

Beginning charge (Before EVA 2): N/A

Ending charge (On return from EVA 2, before recharging): N/A

Currently charging: Yes

Opportunity rover used: Yes

Hours: 76.6

Beginning charge: 100%

Ending charge: 75%

Currently charging: Yes

Curiosity rover used: Yes

Hours: 155.1

Beginning charge: 100%

Ending charge: 97%

Currently charging: Yes

Notes on rovers: One of the rovers was accidentally not turned off after the EVA today. This was identified about three hours after the return from the EVA and resolved with an emergency EVA.

ATV’s Used (Honda, 300, 350.1, 350.2, 350.3): None

Reason for use: N/A

Oil Added? N/A

ATV Fuel Used: N/A

# Hours the ATVs were used today: N/A

Notes on ATVs: N/A

HabCar used and why, where? No

CrewCar used and why, where? No, but we did not report mileage when leaving Grand Junction and heading to MDRS. Mileage – 198640. Tank is full. No damage besides normal wear and tear. Oil was checked — it’s about 75% full (oil level was around the 4th of 5 dots).

General notes and comments: None

Summary of internet: Nominal

Summary of suits and radios: Nominal. The new headset works properly.

Today, the fan on one of the suits slowed down considerably about two hours into the EVA, causing the crew member’s helmet to fog up. The EVA was immediately halted, and the crew attempted to return to the Hab before the fan stopped. Unfortunately, the fan completely failed just about 10 minutes after. It was later found that the suit was only plugged in partially, resulting in an incomplete charge.

Summary of Hab operations: Nominal

Summary of GreenHab operations: Nominal

Summary of ScienceDome operations: Nominal

Summary of RAM operations: Nominal

Summary of any observatory issues: In correspondence with the

astronomy support team.

Summary of health and safety issues: Nominal

Questions, concerns and requests to Mission Support: None

Green Hab Report Dec 23rd

Crew 236 GreenHab Report – 23DEC2021

GreenHab Officer: Vladimir Zeltsman

Environmental Control: Around 17% Humidity

Average Temperatures: 78 degrees

Hours of Supplemental Light: 7 hours

Daily Water Usage For Crops: 9 gallons

Daily Water Usage For Research and/or Other Purposes: None

Water In Blue Tank: 213 Gallons

Times of watering for crops: 8:00 am

Changes to Crops: None

Narrative: None

Harvest: 7g Basil, 20g Mint

Support/supplies needed: None

Astronomy Report Dec 23rd

Crew 236 Astronomy Report 23 Dec 2021

Crew Astronomer:

Name: Dylan Dilger

Crew: 236

Date: 12/23/2021

MDRS ROBOTIC OBSERVATORY

Robotic Telescope Requested: Montana Learning Center

Objects Viewed: None, but processed M37 images taken 2 days ago

Problems Encountered: The same steady wind that affected the Crab nebula images taken the same night elongated the stars

MUSK OBSERVATORY

Solar Features Observed: None

Problems Encountered: Still troubleshooting observatory dome controls, communicating with Astronomy Team about this.

EVA Report Dec 23rd

Crew 236 EVA Report 23DEC2021

EVA #4

Author: Cesare Guariniello

Start time: 9.31 AM

End time: 12.02 PM

Narrative: The team began EVA preparation around 9 and entered the airlock at 9.31, just at the beginning of the EVA window. Crew Journalist replaced his earpiece that caused issues on EVA #3. After depressurization, the crew boarded the vehicles and headed south along Cow Dung road. After a drive of about 5 km, the crew walked towards Barainca Butte. Along the way, XO had a couple of stops to collect rock samples. One rock is brown and black and heavily vesiculated (possibly scoria). Other rocks came down from upper layers of hard mudstone and conglomerates. Meanwhile, SCI took rock temperature measurements for a project on thermal inertia of minerals at MDRS. ENG and JOU assisted and documented the work. The EVA crew reached the south side of Barainca Butte and continued its research proceeding counterclockwise around the butte. After having walked about one quarter of the perimeter of the butte, JOU noted some increasing fogging of the helmet and asked to position the fan switch at maximum speed. The fogging decreased, but JOU reported the fans slowing down. The crew decided to reduce the EVA time and began walking towards the vehicles. Unfortunately, within just a few minutes, JOU experienced a total loss of the fans which, on Mars, would require an urgent contingency rescue or would result in loss of the crewmember. Since the fogging disappeared quickly and JOU reported receiving enough air from underneath the collar, he elected to keep his helmet on for the drive back to the Hab. During the drive back, the crew met some astronauts onboard terrestrial-looking vehicles. The crew stopped the vehicles on the side of Cow Dung road, south of Kissing Camel range, to let the unknown astronauts pass by, with a brief and very amicable exchange of nods and smiles. The crew returned to the Hab about one hour earlier than the expected time, without further incidents.

Destination: Around Barainca Butte

Coordinates (use UTM NAD27 CONUS): N4247200, W519000

Participants: Cesare (XO), Tyler (SCI), Ben (JOU), Pavi (ENG)

Road(s) and routes as per MDRS Map: from the Hab south on Cow Dung Road, west towards Barainca Butte (region J)

Mode of travel: Driving and walking

Vehicles used: Curiosity and Opportunity

Cesare Guariniello

XO/HSO

Crew 236

Journalist Report – December 22nd

Journalist Report
Ben Durkee, Crew 236 Journalist

Sol 03

I did not breach my cocoon like a beautiful butterfly this morning as I had hoped. It was more akin to a greasy, groggy Swamp Thing emerging from a haphazard pile of winter jackets and dirty laundry. But I was warm!

We rose early this morning, in tandem with the sun. Honestly, I don’t know how that thing does it day in and day out. If it weren’t for the dulcet tones of "Ra-Ra Rasputin" pumping us up, I’d probably have stayed in my grease grotto all day.

After wiping the lingering sleep from our eyes and throwing together some cereal, we propelled ourselves into EVA prep.

The plan: myself, Tyler, Cesare, and Pavi were to head farther south than we’ve ever been, past Kissing Camel Ridge all the way to a place called Beranca Butte. Our goal was to find a good spot for Tyler’s thermal analysis research and to survey for any geologically interesting minerals along the way for Cesare. He had his eye set on Hematite.

"But did things go according to plan?" I hear you ask. They hardly ever do. And please don’t interrupt.

We entered the airlock for the first time, began pressurizing, and realized Pavi’s mic wasn’t picking up any audio. An inaudible engineer is no good in a pinch! So, we repressurize, troubleshoot, fix, and enter the airlock once again. A few games of rock, paper, scissors later and we officially made it out of the Habitat.

Two to each rover, we saddled up and hit the road, past mountains and molehills, and Mesozoic masterpieces. We made it about 5 minutes down the road and began experiencing some radio interference so intense that we had to pull over.

After a roadside game of whodunit and a good old-fashioned diagnosis sesh, it turned out my headset had some sort of anomaly that was causing it to constantly transmit, completely gumming up our comms and making communication impossible. In effect, it also gave me an uncontestable soapbox, but I realized too late and I’ll regret that to my dying breath.

We had to turn around and ultimately scrub the EVA due to the time that had elapsed sorting all of this out. Our temporary solution was for me to disconnect my headset and hear their communications by cranking the radio speaker and holding it up to my helmet. It worked! But part of me wishes it hadn’t.

I swear there was a twinkle in Cesare’s eye as it dawned on him that I was alone with the void and had to listen to those knuckleheads the whole ride back and the whole repressurization cycle. A fate I wouldn’t wish upon my worst enemy. Which is now him.

Fortunately, when we returned we were greeted with fresh bread, courtesy of Vladimir! We’ll tackle the same EVA plan tomorrow – this time with far more thorough radio check procedures. We filled the remainder of the EVA slot and then the remainder of the day with personal research time.

For dinner, Tyler and I experimented with combining some fresh herbs, leftover lentils, and dehydrated food into a rice-based amalgamation we called "Jambalaya." It was very much not that. But we all drowned it in seasoning and the crew pretended it was edible, which I appreciated. I foresee a lot more Frankenfoods in the future.

Until then, I’ll recede into my grimy goblin grotto and await more grub.

EVA Report – December 22nd

Crew 236 EVA Report 22-Dec-2021

EVA #3

Author: Cesare Guariniello, XO/HSO Crew 236

Start time: 9.41 AM
End time: 10.28 PM

Narrative: The team first entered the airlock at 9.41. During depressurization, one crew member experienced a problem with the microphone, and after 3 minutes of repressurization the crew came back into the Hab to fix the problem. After the second depressurization, the EVA crew boarded the vehicles and headed south.

About 10 minutes into the drive, loud static was heard in the communication channel. The EVA crew stopped at the side of the road to troubleshoot the problem. XO assessed the radios and found that one radio was stuck in transmission mode (TX), thus preventing one EVA crew member from receiving communication. Switching a radio showed that the issue was in the earpiece, possibly the cable (since the earpiece does not have a vox setting).

As per EVA rules, for reasons of safety, the EVA was aborted due to one crew member has lost communication. The EVA crew returned to the Hab immediately, with no further anomalies.

Destination: Around Baranca Butte (INTENDED)

Coordinates (use UTM NAD27 CONUS): N4247200, W519000

Participants: Cesare (XO), Tyler (SCI), Ben (JOU), Pavi (ENG)

Road(s) and routes as per MDRS Map: from the Hab south on Cow Dung Road, west towards Baranca Butte (region J)

Mode of travel: Driving and walking

Vehicles used: Curiosity and Opportunity

Greenhab Report – December 22nd

GreenHab Officer: Vladimir Zeltsman

Environmental Control: Around 17% Humidity

Average Temperatures: 78 degrees

Hours of Supplemental Light: 7 hours

Daily Water Usage For Crops: 10 gallons

Daily Water Usage For Research and/or Other Purposes: None

Water In Blue Tank: 222 Gallons

Times of watering for crops: 8:00 am

Changes to Crops: None

Narrative: None

Harvest: 17g Green Onion, 3g Parsley, 1g Sage, 1g Thai Hot Peppers

Support/supplies needed: None

Research Report – December 22nd

Crew 236 Astronomy Report Dec 22 2021

Name: Dylan Dilger
Crew: 236
Date: 12/22/2021

MDRS ROBOTIC OBSERVATORY

Robotic Telescope Requested: Montana Learning Center

Objects Viewed: NGC 2244 (Rosette Nebula), also enhanced the M1 crab nebula picture from yesterday now that Lum and Halpha data files have been downloaded. This picture is attached.

Problems Encountered: Is there any update on the MDRS observatory?

MUSK OBSERVATORY

Solar Features Observed: None

Problems Encountered: Still troubleshooting observatory dome controls, communicating with Astronomy Team about this.